OpenStudy (anonymous):

School guidelines require that there must be at least 2 chaperones for every 25 students going on a school trip. How many chaperones must there be for 80 students?

OpenStudy (pixiedust1):

Okay! So, the minimum is 2 per 25. How many times does 25 go into 80?

OpenStudy (anonymous):

3 times

OpenStudy (pixiedust1):

Yes! It goes is evenly 3 times. So 25*3=75 students 2*3=6 chaperones But wait! There's 5 students left over! 25/2=13 (roughly) students. So let's add in an extra chaperone (since you can't have half a human lol)

OpenStudy (pixiedust1):

And a school would rather have more than the state requirement then less

OpenStudy (anonymous):

I got 6 chaperones too, but the answer was wrong.

OpenStudy (pixiedust1):

right, because its 7

OpenStudy (pixiedust1):

Look at what I said. we have to add an extra chaperone to the 6, (which means 6+1=7)

OpenStudy (anonymous):

oh, I got it now. you added they extra human.lol

OpenStudy (pixiedust1):

yeah, because there was 5 extra kids. You need 1 (since 2 is for 25, 1 is for 12. 5 is less then 12, but it's closer than 25)

OpenStudy (pixiedust1):

Make sense?

OpenStudy (anonymous):

I understand now. Thanks. I will award you a medal. :)